OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Pick an option y − 4 = −(x + 6) y − 4 = 2(x + 6) y + 6 = −(x − 4) y + 6 = 2(x − 4)

OpenStudy (anonymous):

I have a 25% chance of being right :D

OpenStudy (anonymous):

m=(0-4)/(2+6) m=1/2 the equation is in form of y2-y1=m(x2-x1) y-4=(1/2)(x+6)

OpenStudy (anonymous):

none are right

OpenStudy (anonymous):

but b looks most promising